# TexShape: Information Theoretic Sentence Embedding for Language Models

### v1

@@ IV-A TexShape to Address Privacy-Utility Trade-Off
In this subsection, we utilize Dataset A and Dataset B. We aim to compress each datasets to filter out sensitive information, while preserving critical information that is needed for certain downstream tasks. Thus, we choose the objective function to be a simplified form of (2), where $\gamma=0$, $N_{l}=N_{s}=1$, $\lambda_{1}=1$ and $\mu_{1}=\mu$, resulting in,
@@
Here, $L(\mathbf{X})$ denotes the “sentiment" label for Dataset A (SST-2) and “location" label for Dataset B (Corona-NLP), and $S(\mathbf{X})$ denotes the “sentence-length" label for Dataset A and “sentiment" label for Dataset B, respectively. Each MI is computed between a compressed representation with size $128$ and a binary label, and thus the network they utilize has $129$ input nodes. Each MI estimation iteration employs a batch size equal to the length of the full data set, since larger batch sizes showed improved numerical stability.
@@
Fig. 2 illustrates the T-SNE mapping obtained before and after employing the TexShape encoder trained with $\mu=0.3$ on Dataset A. The illustrations show that the task-oriented encoder makes distinguishing between samples based on their private label visually impossible, while preserving the visual separability between samples based on their public labels. The estimated MIs between the embedding and the public label is $0.5051$ and $0.4739$ for original and TexShape embeddings. On the other hand, the estimated MIs between the embedding and the private label is $0.5552$ and $0.1334$ for original and TexShape embeddings.
We also considered two classifiers that are trained to predict the public label and the private label of a compressed sample, respectively. We compared three different sentence embedding methods for Dataset B as the training data: Original (when the information-theoretic compression part of the encoder is bypassed), Random (when weights of the the trainable part of the encoder is chosen randomly), and TexShape with $\mu=0.4$. Fig. 3 depicts the predictive performance of these classifiers on unseen data, in terms of their ROC curves, for public label and private label, respectively. As for public label, the area under the ROC curve (AUROC) for TexShape even slightly improves upon the corresponding number for the original representation, suggesting that a compressed version can be more beneficial. As for private label, the AUROC for TexShape is around $0.59$, showing that this embedding can hardly be used by an adversary to identify sensitive information, even when similar public data exists. The performance of random embedding is worse than TexShape, showing lower predictive utility and higher leakage of private information.
@@
Figure 3: Validation ROC Curves for classifiers trained on Dataset B for two different prediction tasks.
@@ IV-B TexShape for Information-Theoretic Compression
In this subsection, we evaluate the compression capability of TexShape without emphasis on or filtering any specific feature. We utilize Dataset C, and we aim to compress textual dataset by transforming the original embedding with size $768$ into a compressed embedding with lower dimensionality, while maximizing the mutual information between the original embedding and the encoded embedding. Thus, the objective function derived from (2) is simplified to $\max_{\Theta}\,\mathcal{I}(T_{\Theta}(\bm{X});\bm{X})$. During TexShape training, only the premise part of each sample was utilized.
For evaluation, we considered the generalization accuracy (ACC.) of two different classifiers, i.e., predictor of the logical relationship between a pair of sentences (Label 1), and predictor of the source of a sentence (Label 2). The classifiers are trained on various embedding of Dataset C and Table I summarizes the results, which manifests the power of our MI-based compression in preserving the predictive utility across these two tasks. We would like to note that training of the encoder was agnostic to any of these features, and regardless, TexShape was successful to preserve notably higher information regarding each of these two features, compared to an untrained (Random) compression.
@@ IV-C TexShape for Compression with Fairness Consideration
In this subsection, we utilize Dataset A and Dataset B, and the focus is to keep maximal information per sample on average, while filtering out certain information that can cause unwanted biases when used for training classifiers. Thus, the objective function is derived from (2) is $\max_{\Theta}\mathcal{I}(T_{\Theta}(\mathbf{X});\mathbf{X})-5\mathcal{I}(T_{\Theta}(\mathbf{X});S(\mathbf{X}))$. Here, $S(X)$ represents the sensitive information in $X$ that needs to be removed, which is the “sentence-length" label for Dataset A and “sentiment" label for Dataset B, respectively. We measure the MI between the embedding and the sensitive label as a measure of bias. Further, to evaluate the utility of the filtered dataset, we evaluated the performance of classifiers that are trained to predict “sentiment" label for Dataset A and “location" label for Dataset B. Table II summarizes the results: As we observe, the measure of bias is dramatically lower in the outputs of TexShape, compared to the original datasets. Such a benefit comes at the cost of a small decrease in the AUROC for a task that TexShape was agnostic to, compared to a similar setting where training was done using original dataset.
@@
Original 768 0.8342 0.5552 0.9566 0.5832
TexShape 128 0.7920 0.1939 0.9284 0.0711
V Conclusions
In this paper, we have integrated fundamental concepts from information theory and machine learning and proposed a semantic approach for sentence embedding in language models. Additionally, we have highlighted applications of our solution in areas such as lossy compression, privacy-preserving data sharing, and training unbiased models. Our work represents an initial step in understanding the relationship between information theory and the rapidly-growing field of large language models.
@@ Acknowledgment
Donsker-Varadhan Representation for MI
Let consider an arbitrary well-behaved function $F:\Omega\rightarrow\mathbb{R}$. We define another distribution $G:\Omega\rightarrow\mathbb{R}$ (called Boltzmann distribution in the literature) as follows:
$G(X,Z)=\frac{e^{F(X,Z)}}{E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}]}P_{\bm{X}}(X)P_{\bm{Z}}(Z).$
Because of linearity of expectation, we have,
$-\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}]){=}\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)}){-F}(X,Z).$ (3)
Since the left hand side of (3) is deterministic, the right hand side must also be deterministic, thus the expectation does not change its value,
$-\log(E$ [${}_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])=$
$E_{P_{\bm{X,Z}}}[\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})]-E_{P_{\bm{X,Z}}}[F(X,Z)].$
Leading to,
$E_{P_{\bm{X,Z}}}[F(X,Z)]-\log(E$ [${}_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])=$
$E_{P_{\bm{X,Z}}}[\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})].$
On the other hand, by definition of mutual information, we have
$\mathcal{I}(\bm{X};\bm{Z})=E_{P_{\bm{X,Z}}}[\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})].$ (4)
Therefore,
$\begin{split}\Delta&=\mathcal{I}(\bm{X};\bm{Z})-\left[E_{P_{\bm{X,Z}}}[F(X,Z)]-\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])\right]\\ &=E_{P_{\bm{X,Z}}}[\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})-\log(\frac{G(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})]\\ &=E_{P_{\bm{X,Z}}}[\log(\frac{P_{\bm{X,Z}}(X,Z)}{G(X,Z)})]\\ &=D_{\text{KL}}(P_{\bm{X,Z}}(X,Z)||G(X,Z))\geq 0\end{split}$
As a result,
$E_{P_{\bm{X,Z}}}[F(X,Z)]-\log(E_{P_{\bm{X}}P_{\bm{Z}}}[e^{F(X,Z)}])\leq\mathcal{I}(\bm{X};\bm{Z}).$
The bound is achieved when $D_{\text{KL}}(P_{\bm{X,Z}}(X,Z)||G(X,Z))=0$. This happens only when $G(X,Z)=P_{\bm{X,Z}}(X,Z)$. Equivalently, when
$F(X,Z)=\log(\frac{P_{\bm{X,Z}}(X,Z)}{P_{\bm{X}}(X)P_{\bm{Z}}(Z)})+C,$ (5)
for a constant $C$.

### v2

@@ IV-A TexShape to Address Privacy-Utility Trade-Off
In this subsection, we utilize Dataset A and Dataset B. We aim to compress each datasets to filter out sensitive information while preserving critical information that is needed for certain downstream tasks. Thus, we choose the objective function to be a simplified form of (2), where $\gamma=0$, $N_{l}=N_{s}=1$, $\lambda_{1}=1$ and $\mu_{1}=\mu$, resulting in
@@
Here, $L(\mathbf{X})$ denotes the “sentiment" label for Dataset A (SST-2) and “location" label for Dataset B (Corona-NLP), and $S(\mathbf{X})$ denotes the “sentence-length" label for Dataset A and “sentiment" label for Dataset B, respectively. Each MI is computed between a compressed representation with size $128$ and a binary label, and thus the network they use has $129$ input nodes. Each MI estimation iteration employs a batch size equal to the length of the full data set.
@@
Fig. 2 illustrates the T-SNE mapping obtained before and after employing the TexShape encoder trained with $\mu=0.3$ on Dataset A. The illustrations show that the task-oriented encoder makes distinguishing between samples based on their private label visually impossible, while preserving the visual separability between samples based on their public labels. The estimated MIs between the embedding and the public label are $0.5051$ and $0.4739$ for original and TexShape embeddings, respectively. On the other hand, the estimated MIs between the embedding and the private label are $0.5552$ and $0.1334$ for the original and TexShape embeddings, respectively.
We also considered two classifiers trained to predict the public label and the private label of a compressed sample, respectively. We compared four different sentence embedding methods for Dataset B as the training data: Original (when the information-theoretic compression part of the encoder is bypassed), Random (when weights of the trainable part of the encoder are chosen randomly), Noisy (wherein the pixels of training data undergo independent distortion with Gaussian additive noise of mean 00 and variance $\sigma^{2}$, aligned with the perspective of differential privacy), and TexShape with $\mu=0.4$. Fig. 3 depicts the predictive performance of these classifiers on unseen data, in terms of their ROC curves, for public label and private label, respectively. As for the public label, the area under the ROC curve (AUROC) for TexShape even slightly improves the corresponding number for the original representation, suggesting that a compressed version can be more beneficial. As for private label, the AUROC for TexShape is around $0.51$, showing that this embedding can hardly be used by an adversary to identify sensitive information, even when similar public data exist. The performance of random embedding and noisy embedding is worse than TexShape, showing lower predictive utility and higher leakage of private information.
@@
Figure 3: Validation ROC and AUROC for classifiers trained on Dataset B for public task (left) and private task (right).
@@ IV-B TexShape for Information-Theoretic Compression
In this subsection, we evaluate the compression capability of TexShape without emphasizing or filtering any specific feature. We utilize Dataset C, and we aim to compress textual dataset by transforming the original embedding with size $768$ into a compressed embedding with lower dimensionality, while maximizing mutual information between the original embedding and the encoded embedding. Thus, the objective function derived from (2) is simplified to $\max_{\Theta}\,\mathcal{I}(T_{\Theta}(\bm{X});\bm{X})$. During TexShape training, only the premise part of each sample was utilized.
For evaluation, we considered the generalization accuracy (ACC.) of two different classifiers, that is, predictor of the logical relationship between a pair of sentences (Label 1) and predictor of the source of a sentence (Label 2). The classifiers are trained on various embeddings of Dataset C and Table I summarizes the results, which manifests the power of our MI-based compression in preserving predictive utility in these two tasks. We would like to note that encoder training was agnostic to any of these features, and regardless, TexShape was successful in preserving higher information regarding each of these two features compared to untrained compression.
@@ IV-C TexShape for Compression with Fairness Consideration
In this subsection, we utilize Dataset A and Dataset B, and the focus is on keeping maximal information per sample on average, while filtering out certain information that can cause unwanted biases. Thus, the objective function derived from (2) is $\max_{\Theta}\mathcal{I}(T_{\Theta}(\mathbf{X});\mathbf{X})-5\mathcal{I}(T_{\Theta}(\mathbf{X});S(\mathbf{X}))$. Here, $S(X)$ represents the “sentence-length" label for Dataset A and “sentiment" label for Dataset B, respectively. We measure the MI between the embedding and the sensitive label as a measure of bias. Further, to evaluate the utility of the filtered dataset, we evaluated the performance of classifiers that are trained to predict “sentiment" label for Dataset A and “location" label for Dataset B. Table II summarizes the results: As we observe, the bias measure is dramatically lower in the outputs of TexShape, compared to the original datasets. This benefit comes at the cost of a small decrease in AUROC for a task that TexShape was agnostic to, compared to a similar setting where training was performed using the original dataset. Our fairness method here is conceptually aligned with [30] where compression is used to de-bias the GloVe embedding. However, we do not restrict ourselves to a specific embedding.
@@
Original 768 0.9566 0.5832 0.8443 0.5552
TexShape 128 0.9284 0.1939 0.7920 0.0711
V Discussion and Future Work
In this paper, we have integrated fundamental concepts from information theory and machine learning and proposed a semantic approach for data processing in language models. We then highlighted the applications of our solution in areas such as lossy compression, privacy-preserving data sharing, and training unbiased models. Our proposed design objective which is based on a weighted linear combination of data processing goals such as resource utilization, task-specific utility, privacy, and fairness, stands as a foundation for further exploration. For instance, expanding the linear combination to more general forms is promising in achieving a better balance among competing goals. Furthermore, optimizing the weights as hyperparameters, informed by theory, can bridge a link between design choices and desired performance metrics. These avenues are key directions for our future studies.